# ConCodeEval: Evaluating Large Language Models for Code Constraints in Domain-Specific Languages

### v2

@@ A.1.3 Llama family
Other than this, similar to the granite family we try Question answering format and instruction to wrap the output in quotes (“‘).
Output Representation
JSON YAML XML
Model Schema Gen Acc Val Acc Gen Acc Val Acc Gen Acc Val Acc
Llama3 8B JSON 48.3 71.2 46.6 68.1 39.2 64.1
Granite 8B 51.2 69.2 52.3 66.1 47.8 65.8
Granite 20B 58.3 73.5 56.4 72.3 50.2 68.2
Granite 34B 66.3 76.2 64.5 75.4 51.3 73.2
Codellama 34B 65.1 75.1 63.4 73.2 50.6 71.2
[…]
Llama3 8B XML 46.6 65.8 42.3 63.4 36.6 60.1
Granite 8B 46.2 64.8 44.5 63.2 34.5 57.3
Granite 20B 50.4 66.7 48.2 64.1 36.4 56.1
Granite 34B 52.3 68.5 51.1 63.4 39.2 53.2
Codellama 34B 49.2 66.2 49.2 63.2 35.1 52.1
[…]
Llama3 8B YAML 46.7 67.2 45.3 64.2 43.5 63.2
Granite 8B 48.1 65.2 46.2 61.2 44.2 61.2
Granite 20B 52.3 68.9 49.7 66.7 47.8 65.1
Granite 34B 54.2 67.7 51.3 65.3 45.3 56.4
Codellama 34B 56.8 66.4 50.2 64.3 47.8 56.2
[…]
Llama3 8B Python 43.2 60.1 41.1 58.9 39.2 57.6
Granite 8B 45.1 60.5 46.7 59.4 37.4 56.0
Granite 20B 48.2 57.2 45.9 57.8 38.4 58.2
Granite 34B 50.6 59.2 47.1 55.6 41.3 57.3
Codellama 34B 47.2 56.4 45.3 57.2 39.2 55.1
[…]
A.2 Data statistics
This section represents schema length comparison for various languages.
Language Max schema length Average schema length
XML $3316$ $364.82$
JSON $1954$ $208.23$
YAML $1295$ $135.09$
Table 3: Schema length comparison
Constraint Llama3 8B Llama3 70B
type 302 49
exclusiveMinimum 18 44
multipleOf 170 42
minLength 47 21
contains 22 12
exclusiveMaximum 22 12
maximum 11 2
maxLength 7 19
additionalProperties 4 0
minimum 4 15
Table 4: Both the models, least and best performing irrespective of their performance show a similar distribution of mistakes for each constraint.
A.3 Few shot prompting results
Below are the results for Few-shot prompting. We experiment with 3 shot prompting. We observe that the majority of errors made by all the models are regarding short schema and the schema having root type of array as shown in sample LABEL:lst:task1. An example of a 3-shot prompt for a DSL generation task is shown below.
@@ A.3 Few shot prompting results
A.4 Limitations of Constrained Decoding
This section outlines some common problems with constrained decoding and emphasises on why it cannot be a complete and a viable solution for factoring in schemas to generate compliant text using language models.
A.4.1 Inference Performance Bottleneck
Constrained decoding often negatively effects inference throughput widely mentioned as one of the major drawbacks in many works Wang et al. (2024b); Pimparkhede et al. (2024); Geng et al. (2023) due to involvement of token-level operations keeping track of the schema constraints and tokens generated so far. This latency can be a factor of the complexity of the schema, tokens generated so far, and the nature of the constrained decoding implementation. Further, advances such as batched inference  [^6] https://github.com/microsoft/batch-inference are not yet there for constrained decoding limiting their scalability and practical use.
A.4.2 Complex Engineering Effort
@@
A.4.3 Model Performance Bottleneck
@@ A.4.4 Limited Scope
A.5 Task Motivation
A.5.1 Data as Code Generation Task
This section describes use cases from enterprise and research point of view motivating data as code generation seed task in our study.
@@ A.5.1 Data as Code Generation Task
(i) Test case structured data generation to test application interfaces such as REST API endpoints. Often enterprises have large number of services exposing API endpoints that have to be tested and LLMs can be a drop in solution to generate test case data at scale. (ii) Structured configuration data generation for a particuilar use case and domain. Enterprise applications such as Kubernetes use DSLs for configuration and usage, preparing them require deep domain expertise and there is increasing motivation Pujar et al. (2023) to employ LLMs in enterprises to generate DSL code. (iii) Some more downstream tasks involving structured data such as forms and tables often represented in programmable format such as JSON can leverage LLMs to generate structured data to fill forms or tables leveraging the schema.
@@
A.5.2 DSL Validation Task
This section describes use cases from enterprise and research point of view motivating DSL validation seed task in our study.
@@ A.5.2 DSL Validation Task
(i) Given the schema, employing LLMs to generate domain aware suggestions over the provided structured data which is not viable with traditional schema validators which only pinpoint syntactic errors and cannot provide semantic suggestions. Such as providing optimizations over the existing resource YAML in Kubernetes while complying with resource schema. (ii) In a assistive chat system, often the constraints are in NL representation from the user which are not machine readable and LLMs should be able to understand such constraints. (iii) Quick interoperability across different schema and data representation versions. Often in enterprises, schemas can be in a particular version not compatible with structured data’s version. For instance, the schema could be in an older JSON schema version such as Draft 0 and data in Draft 7, in such cases LLMs can come handy to perform validation at scale.
@@
Understanding LLMs’ capability in validating the given structured data against the schema across representations can provide seed evidence for more complex tasks such as automatic fixing of data in compliance with given schema.
A.6 Schema Examples
@@
A.7 Hyperparameter details
We perform inference for all the models in $float16$ precision and a max new token limit of 1024 tokens. For beam search decoding, we use the beam width of 3.
A.8 Data Generation Scripts
A combinatorial data generation tool is created, which factors in constraints of interest, constraint-specific information, and combinatorial preferences, to generate the schemas. We also use openly available automatic lossless language to language translation tools for translation code from JSON to YAML etc. We plan to open-source all the scripts used for data preparation.

### v3

@@ A.1.3 Llama family
Other than this, similar to the granite family, we try Question answering format and instruction to wrap the output in quotes (“‘).
@@
Figure 2: Attention maps for LLama3 8B and 70B model for Data validation experiment. The more the intensity of color, the more attention is given to that part of input by the model.
A.2 Limitations of Constrained Decoding
This section outlines some common problems with constrained decoding and emphasizes why it cannot be a complete and viable solution for factoring in schemas to generate compliant text using language models.
A.2.1 Inference Performance Bottleneck
Constrained decoding often negatively affects inference throughput, widely mentioned as one of the major drawbacks in many works Wang et al. (2024); Pimparkhede et al. (2024); Geng et al. (2023) due to involvement of token-level operations keeping track of the schema constraints and tokens generated so far. This latency can be a factor of the complexity of the schema, tokens generated so far, and the nature of the constrained decoding implementation. Further, advances such as batched inference  [^5] https://github.com/microsoft/batch-inference are not yet there for constrained decoding limiting their scalability and practical use.
A.2.2 Complex Engineering Effort
@@
Constraint Llama3 8B Llama3 70B
type 302 49
exclusiveMinimum 18 44
multipleOf 170 42
minLength 47 21
contains 22 12
exclusiveMaximum 22 12
maximum 11 2
maxLength 7 19
additionalProperties 4 0
minimum 4 15
Table 3: Both the models, least and best performing, irrespective of their performance, show a similar distribution of mistakes for each constraint.
A.2.3 Model Performance Bottleneck
@@ A.2.4 Limited Scope
Language Max schema tokens Avg schema tokens
XML $3316$ $364.82$
JSON $1954$ $208.23$
YAML $1295$ $135.09$
Table 5: Schema length comparison using Llama3 tokenizer
A.3 Task Motivation
A.3.1 Data as Code Generation Task
This section describes use cases from enterprise and research points of view motivating data as code generation seed tasks in our study.
@@ A.3.1 Data as Code Generation Task
(i) Test case structured data generation to test application interfaces such as REST API endpoints. Often, enterprises have a large number of services exposing API endpoints that have to be tested, and LLMs can be a drop-in solution to generate test case data at scale. (ii) Structured configuration data generation for a particular use case and domain. Enterprise applications such as Kubernetes use DSLs for configuration and usage, preparing them require deep domain expertise and there is increasing motivation Pujar et al. (2023) to employ LLMs in enterprises to generate DSL code. (iii) Some more downstream tasks involving structured data, such as forms and tables often represented in a programmable format such as JSON, can leverage LLMs to generate structured data to fill forms or tables leveraging the schema.
@@
A.3.2 DSL Validation Task
This section describes use cases from an enterprise and research perspective that motivate our study’s DSL validation seed task.
@@ A.3.2 DSL Validation Task
(i) Given the schema, employing LLMs to generate domain-aware suggestions over the provided structured data is not viable with traditional schema validators, which only pinpoint syntactic errors and cannot provide semantic suggestions. Such as providing optimizations over the existing resource YAML in Kubernetes while complying with resource schema. (ii) In an assistive chat system, the constraints are often in NL representation from the user, which is not machine-readable, and LLMs should be able to understand such constraints. (iii) Quick interoperability across different schema and data representation versions. Often in enterprises, schemas can be in a particular version that is incompatible with the structured data version. For instance, the schema could be in an older JSON schema version such as Draft 0 and data in Draft 7, in such cases LLMs can come handy to perform validation at scale.
@@
Understanding LLMs’ capability in validating the given structured data against the schema across representations can provide seed evidence for more complex tasks such as automatically fixing data in compliance with the given schema.
A.4 Schema Examples